1. log base 5 of 25radical5

Answer by grishma.kshatriya(27)

Solution :
log to the base 5 (25)^5
Applying the rule as log(b)^a = a log(b)
5 log to the base 5 (25)
(25) = (5)^2
So
5 log to the base 5 (5)^2
Also the rule says that log to the base a (a)^b = b
5 * 2
= 10
Thank you .